OpenStudy (anonymous):

One glucosamine pill has a mass of 1.14g. How much will a pharmacist weigh out to fill a prescription that calls for 500 pills. I got 570 g B) how many glucosamine pills are there in 88kg? Not sure how to solve this

OpenStudy (aaronq):

for both scenarios, you need a use a ratio. we know, \(\sf 1.14g=~1pill\rightarrow \dfrac{1.14~g}{1~pill}\) \(\sf \dfrac{1.14~g}{1~pill}=\dfrac{x}{500~pills}\) solve for x, the mass of 500 pills. b) \(\sf \dfrac{1.14~g}{1~pill}=\dfrac{88~kg}{x}\) now before we go ahead and solve this, notice that the units of mass are different. We need to convert 88 kg to g before carrying on. \(\sf 1~kg=1000~g \rightarrow \dfrac{1000~g}{1~kg}\) \(\sf 88~\cancel{kg}* \dfrac{1000~g}{1~\cancel{kg}}=88000~g\) note the units cancel out. so, it looks like \(\sf \dfrac{1.14~g}{1~pill}=\dfrac{88000~g}{x}\) solve for x, the number of pills.